Express the radius of the Earth in unit of solar radius. (Radius of the earth is 6.4 x 10^6m).
1
Expert's answer
2019-05-29T09:29:43-0400

Let's express the radius of the Earth in unit of solar radius (1 solar radius is equal to "6.96 \cdot 10^8 m"):


"R_{Earth} = 6.4 \cdot 10^6 m = 6.4 \cdot 10^6 m \cdot (\dfrac{1R_{\odot}}{6.96 \cdot 10^8 m }) = 0.009 R_{\odot}."

Answer:

"R_{Earth} = 0.009 R_{\odot}."